OpenStudy (anonymous):

calculate the number of atoms in 168.0 HgS - can someone please show me how to do this with the steps, I got a answer, but im not sure if its correct

OpenStudy (aaronq):

you need to find the molar mass of the compound, look at the periodic table and add up Hg + S. then you divide the grams by the molar mass of the compound \[moles =\frac{ mass (g) }{molar mass(g/mol) }\]

OpenStudy (anonymous):

i got the final answer as 8.7x10^23 is tht correct? in the end do yu have to multiply the answer with the number of atoms yu have? because thats what i did

OpenStudy (aaronq):

sorry i forgot to mention that, you do multiply the number of moles by avogadros number.. 6.023x10^23

OpenStudy (aaronq):

but your number seems high because 168 g is not even one mole

OpenStudy (anonymous):

so what is the answer then? can yu calculate it and tell me pls

OpenStudy (aaronq):

moles = [168/(200.59+32.066)] atoms = (moles) (6.023x10^23) =434918506292552094078811.636063544460491025376518734

OpenStudy (anonymous):

i got tht number first but then i multiplied it by 2 since there are 2 atoms in it

OpenStudy (anonymous):

HgS <-- 2 atoms?

OpenStudy (anonymous):

isnt that how yu do it

OpenStudy (aaronq):

haha frig, you're so right, idk why I'm not thinking

OpenStudy (anonymous):

LOOL

OpenStudy (anonymous):

yay i am right, so if this question was asking for the number of molecules, would you still have to multiply the number of atoms or no?

OpenStudy (aaronq):

if it asked for the number of molecules, then you don't multiply by 2

OpenStudy (anonymous):

oh arii thanks !

OpenStudy (aaronq):

no probs !